# Dream-in-Style: Text-to-3D Generation using Stylized Score Distillation

### v1

@@ 4.3. Optimization
(13) $\tau_{\text{sqrt}}(\lambda;\lambda_{\text{max}},k,K)=\lambda_{\text{max}}\sqrt{\frac{k}{K}},$
@@
(14) $\tau_{\text{quad}}(\lambda;\lambda_{\text{max}},k,K)=\lambda_{\text{max}}\left(\frac{k}{K}\right)^{2},$
@@ 5. Experimental Results
We perform several experiments to demonstrate the effectiveness of our proposed method. First, we compare our method with three baseline methods for stylized text-to-3D generation. Second, we demonstrate that our method can be adapted to other score distillation losses. Finally, we provide ablation studies to validate the importance of our combined score distillation, as well as perform parameter studies to validate our style ratio scheduling. We perform quantitative evaluation of our method through a human-like user study using large language models (Wu et al., 2024).
Figure 2. Ablation studies. We confirm the effectiveness of our combined distillation by comparing with two baselines: A) the vanilla text-to-3D generation without styles, B) text-to-3D generation guided by only the modified pretrained model.
5.1. Implementation details
@@
We use the implementation of neural radiance fields from threestudio (Guo et al., 2023), which is based on NerfAcc (Li et al., 2023a), as the 3D representation for our optimization. We follow the implementation of DreamFusion (Poole et al., 2023), noise-free score distillation (Katzir et al., 2024) and ProlificDreamer (Wang et al., 2023b) to implement the score distillations. In our method, we apply an augmentation of the text prompt $y$ by concatenating it with a BLIP2 generated caption (Li et al., 2023b) of the style reference image, and use this augmented prompt for the modified score $\hat{\epsilon}$. We set the classifier-free guidance (CFG) scale to 100 for score distillation sampling (SDS) (Poole et al., 2023), and 7.5 for noise-free score distillation (NFSD) (Katzir et al., 2024), and variational score dis- tillation (VSD) (Wang et al., 2023b). We use NFSD as the default score distillation for our method.
@@
Our experiments are performed on a NVIDIA RTX 4090 GPU with 24 GB of VRAM. Our method optimized a 3D object in approximately 1.5 hours using SDS (Poole et al., 2023) or NFSD (Katzir et al., 2024) and 2.5 hours using VSD (Wang et al., 2023b), similar to the training time of the vanilla implementations of these methods.
5.2. Qualitative results
@@
We compare our method with the baseline methods proposed in Section 3. Figure 5 presents a list of text prompts and style reference images with the corresponding outputs of all methods. As can be seen, our results have the best visual quality with consistent alignment to the input pairs of text prompts and reference images.
@@
To demonstrate the adaptability of our method to other score distillation losses, we apply our method to the vanilla score distillation sampling (SDS), noise-free score distillation (NFSD), and variational score distillation (VSD). Figure 6 provides the results of our method adapted to these score distillation losses. It can be seen that the NFSD variant works best, outperforming SDS and VSD. Our method, however, does not yet adapt the LoRA model for scoring noisy rendering in VSD for stylization. Incorporating LoRA with attention feature swapping for stylization might improve the performance of VSD, which is the subject of future work.
@@
Figure 3. Effects of our schedule function with different style ratios on the stylization results. We found that using $sqrt$ schedule with $\lambda=0.6$ works well in general.
5.3. Ablation studies
@@
Figure 4. The results of GPTEval3D on text-geometry alignment, text-asset alignment, style alignment, geometry details, texture details, and 3D plausibility. Our method is effective on all evaluation aspects.
5.4. Quantitative results
We aim to conduct a user study to quantitatively evaluate our method. The existence of large language models with vision capability (e.g., GPT-4v model (OpenAI, 2023b)) allows us to prompt a language model for 3D asset evaluation, which had been demonstrated to work well for the text-to-3D generation task with human-like performance (Wu et al., 2024). We follow (Wu et al., 2024) and extend their GPTEval3D tool to incorporate style evaluation, resulting to six evaluation criteria including text-geometry alignment, text-asset alignment, style alignment, geometry details, texture details, and 3D plausibility. We use this tool to compare the results generated by four methods, including style-in-prompt, neural style loss, textual inversion, and our method. We set up the style-in-prompt method as the base/anchor for the evaluation process. We asked GPTEval3D to perform 120 pairwise comparisons, and then calculated the Elo score (Elo, 1966) for each method. The results are presented in Figure 4. It can be seen that our method outperforms all baseline methods in this evaluation.
6. Conclusions and limitations
@@
Our method is not without limitations. A particular problem is that our results are prone to the Janus problem (Poole et al., 2023), which could be mitigated by using a pretrained diffusion model trained on multi-view datasets. Additionally, our method is best compatible with SDS-family losses. Adapting our method to non-SDS losses is future work. It is of great interest to extend the investigation of stylized text-to-3D generation to videos and 4D data for style references and outputs, respectively.
Appendix A Derivation
Recall that our stylized score distillation aims to minimize $KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))$, which is equivalent to
(15) $\min_{\theta}\mathbb{E}_{\varepsilon}\left[\log(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta)))-(1-\lambda)\log(p_{\phi}(\mathbf{z}_{t}\mid y))-\lambda\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right].$
@@ Appendix A Derivation
(16) $\nabla_{\theta}KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))=\mathbb{E}_{\varepsilon}\left[\underbrace{\nabla_{\theta}\log(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta)))}_{(A)}-(1-\lambda)\underbrace{\nabla_{\theta}\log(p_{\phi}(\mathbf{z}_{t}\mid y))}_{(B)}-\lambda\underbrace{\nabla_{\theta}\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))}_{(C)}\right].$
@@
(17) $\nabla_{\theta}\log(p_{\phi}(\mathbf{z}_{t}\mid y))=-\frac{\alpha_{t}}{\sigma_{t}}{\varepsilon_{\phi}}(\mathbf{z}_{t}\mid y)\frac{\partial{\mathbf{x}}}{\partial\theta},$
@@
(18) $\nabla_{\theta}\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))=-\frac{\alpha_{t}}{\sigma_{t}}\hat{\varepsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})\frac{\partial{\mathbf{x}}}{\partial\theta}.$
@@
(19) $\nabla_{\theta}\log q(\mathbf{z}_{t}|\mathbf{x})=\left(\underbrace{\frac{\partial\log q(\mathbf{z}_{t}|\mathbf{x})}{\partial x}}_{\text{parameter score}}+\underbrace{\frac{\partial\log q(\mathbf{z}_{t}|\mathbf{x})}{\partial\mathbf{z}_{t}}\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{x}}}_{\text{path derivative}}\right)\alpha_{t}\frac{\partial\mathbf{x}}{\partial\theta}=\left(\frac{\alpha_{t}}{\sigma_{t}}{\epsilon}-\frac{\alpha_{t}}{\sigma_{t}}{\epsilon}\right)\alpha_{t}\frac{\partial\mathbf{x}}{\partial\theta}=0.$
Similarly to DreamFusion (Poole et al., 2023), based on Sticking-the-Landing (Roeder et al., 2017), we can discard the parameter score term and keep only the path derivative term.
@@
$\nabla_{\theta}\mathcal{L}_{SSD}$ $=\mathbb{E}_{t,\mathbf{z}_{t}\mid\mathbf{x}}\left[\omega(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right]$
(20) $=\mathbb{E}_{t,\varepsilon}\left[\omega(t)\left((1-\lambda)\varepsilon_{\phi}(\mathbf{z}_{t}\mid y)+\lambda\hat{\varepsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\varepsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right].$
Appendix B GPT-4 Evaluation Template
We extend the GPTEval3D toolbox (Wu et al., 2024) to evaluate the style alignment between the generated 3D objects and the style reference image. The toolbox prompts the GPT-4 language model from OpenAI to perform the evaluation. The entire text prompt to GPT-4 is listed below. Instruction #6 is for style alignment evaluation. Figure 7 shows an example of the image grid sent to GPT-4 for evaluation. In Table 1, we present the detailed Elo scores obtained from this evaluation.
@@ 7. Overall, xxxxxx
Table 1. Detailed results from GPTEval3D evaluation.
Figure 7. An image grid sent to GPT for evaluation. The first and second column show the color rendering and normal map of the first method, and the third and fourth column are for the second method. Each row shows a camera view of the objects, and the last row shows the style reference. The prompt is ”a toy car” and the style image represents fire on a black background. GPT is asked to pick the better result out of the two presented methods.

### v2

@@ 4.3 Optimization
$\tau_{\text{sqrt}}(\lambda;\lambda_{\text{max}},k,K)=\lambda_{\text{max}}\sqrt{\frac{k}{K}},$ (13)
@@
$\tau_{\text{quad}}(\lambda;\lambda_{\text{max}},k,K)=\lambda_{\text{max}}\left(\frac{k}{K}\right)^{2},$ (14)
@@ 5 Experimental Results
We perform several experiments to demonstrate the effectiveness of our proposed method. First, we compare our method with three baseline methods for stylized text-to-3D generation. Second, we demonstrate that our method can be adapted to other score distillation losses. Finally, we provide ablation studies to validate the importance of our combined score distillation, as well as perform parameter studies to validate our style ratio scheduling. We perform quantitative evaluation of our method through a human-like user study using large language models [74].
Figure 2: Qualitative comparisons to baseline methods. Our stylized score distillation leads to 3D generation results consistently aligned with the text prompts and the style reference images. Multiple view rendering are shown in the supplementary video.
Figure 3: Our method applied to different score distillation losses. Multiple view rendering are shown in the supplementary video.
5.1 Implementation details
@@
We use the implementation of neural radiance fields from threestudio [19], which is based on NerfAcc [39], as the 3D representation for our optimization. We follow the implementation of DreamFusion [59], noise-free score distillation [34] and ProlificDreamer [72] to implement the score distillations. In our method, we apply an augmentation of the text prompt $y$ by concatenating it with a BLIP2 generated caption [38] of the style reference image, and use this augmented prompt for the modified score $\hat{\epsilon}$. We set the classifier-free guidance (CFG) scale to 100 for score distillation sampling (SDS) [59], and 7.5 for noise-free score distillation (NFSD) [34], and variational score distillation (VSD) [72]. We use NFSD as the default score distillation for our method.
@@
Our experiments are performed on a NVIDIA RTX 4090 GPU with 24 GB of VRAM. Our method optimized a 3D object in approximately 1.5 hours using SDS [59] or NFSD [34] and 2.5 hours using VSD [72], similar to the training time of the vanilla implementations of these methods.
5.2 Qualitative results
@@
We compare our method with the baseline methods proposed in Section 3.2. Figure 2 presents a list of text prompts and style reference images with the corresponding outputs of all methods. As can be seen, our results have the best visual quality with consistent alignment to the input pairs of text prompts and reference images.
Figure 4: Our results on various style images from  [22].
Figure 5: Our results on complex and detailed text prompts.
@@
To demonstrate the adaptability of our method to other score distillation losses, we apply our method to the vanilla score distillation sampling (SDS), noise-free score distillation (NFSD), and variational score distillation (VSD). Figure 3 provides the results of our method adapted to these score distillation losses. It can be seen that the NFSD variant works best, outperforming SDS and VSD. Our method, however, does not yet adapt the LoRA model for scoring noisy rendering in VSD for stylization which is the subject of future work.
To demonstrate the robustness of our method across different styles and complex prompts, we tested it on various styles as in [22], as shown in Figure 4. Additionally, we applied our approach to more complex textual prompts, which are illustrated in Figure 5.
To verify the robustness of our stylized score distillation across training-free methods for style transfer, we attempted our method on StyleAligned [22]. Compared to visual style prompting (VSP) [30] that only attends to the style image, StyleAligned attends to both the style image and the resulting image. We observed that our SSD with VSP has more natural results. For example, given a fire style image and prompt ’a toy car’, SSD with StyleAligned generates fire around a car, while VSP produces a car made of fire (similar trend with their 2D results). A more detailed discussion is provided in the supplementary material.
5.3 Ablation studies
Figure 6: Ablation studies. We confirm the effectiveness of our combined distillation by comparing with two baselines: A) the vanilla text-to-3D generation without styles, B) text-to-3D generation guided by only the modified pretrained model.
@@ 5.3 Ablation studies
Figure 7: Effects of our schedule function with different style ratios on the stylization results.
@@
Figure 8: Results of GPTEval3D on text-geometry alignment, text-asset alignment, style alignment, geometry details, texture details, and 3D plausibility, which confirm the effectiveness of our method.
5.4 Quantitative results
The existence of large language models with vision capability (e.g., GPT-4v model [54]) allows us to prompt a language model for 3D asset evaluation, which had been demonstrated to work well for the text-to-3D generation task with human-like performance [74]. We follow [74] and extend their GPTEval3D tool to incorporate style evaluation, resulting to six evaluation criteria including text-geometry alignment, text-asset alignment, style alignment, geometry details, texture details, and 3D plausibility. We use this tool to compare the results generated by four methods, including style-in-prompt, neural style loss, textual inversion, and our method. We set up the style-in-prompt method as the base/anchor for the evaluation process. We asked GPTEval3D to perform 120 pairwise comparisons, and then calculated the Elo score [12] for each method. The results are presented in Figure 8. It can be seen that our method outperforms all baseline methods in this evaluation.
6 Conclusions and limitations
@@
Our method is not without limitations. A particular problem is that our results are prone to the Janus problem [59], which could be mitigated by using pretrained diffusion models for multi-view generation [63, 71]. Additionally, our method is best compatible with SDS-family losses. Adapting our method to non-SDS losses is future work. It is of great interest to extend the investigation of stylized text-to-3D generation to videos and 4D data for style references and outputs, respectively.
7 Acknowledgments
This work was conducted with the financial support of the Research Ireland Centre for Research Training in Digitally-Enhanced Reality (d-real) under Grant No. 18/CRT/6224. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
This project is supported by Research Ireland under the Research Ireland Frontiers for the Future Programme, award number 22/FFP-P/11522.
Supplementary Material
In this supplementary document, we (1) further discuss our limitations and potential mitigations of the Janus problem using MVDream. We then (2) provide more evaluations of our method and compare it with image-to-3D methods as well as a second style transfer method. We also (3) provide a detailed derivation of our stylized score distillation, and (4) more details of our GPT-based user study. For more qualitative results, please see our supplementary video.
A Further discussion on limitations
A.1 Multi-face Janus problem
Our method inherits the Janus problem [59] known to score distillation sampling because we use Stable Diffusion, a single-view image diffusion model as the 3D generation prior. To mitigate this problem, one can perform score distillation with multi-view diffusion models. Here, we demonstrate that our stylized score distillation (SSD) extends naturally to MVDream, a popular multi-view diffusion model [63].
In Figure 9 we present cases where multi-face Janus problem is mitigated when our method is combined with MVDream. Multi-view videos of these examples can be found in our supplementary video.
Figure 9: Example cases where multi-face Janus problem occurred and was mitigated by using MVDream.
A.2 Discussion on hard cases
We found style images with ambiguous content or complex backgrounds particularly challenging, e.g., fire style from Figure 9 or cloud style from Figure 10. We found that our style ratio scheduling is particularly important to address these cases during optimization. Without proper scheduling, these cases produced poor results with little or no information about the object in the final results.
B More evaluations
B.1 Image-to-3D on stylized images
We experimented with lifting stylized images to 3D using CRM and TripoSR, two popular large reconstruction models for image-to-3D reconstruction. CRM also uses ImageDream [71], an image-conditioned multi-view diffusion model as an initialization to train their reconstruction model. We found that there are two limitations to this approach. First, when the stylized images have ambiguous 2D geometry (fire around a car), the 3D results by CRM and TripoSR are worse than ours. Second, CRM and TripoSR are trained with clean object rendering, which does not generalize well to stylized images with complex visual effects. Comparison of these cases can be found in Figure 10.
Figure 10: Comparison of our method with results from image-to-3D on stylized images.
B.2 Visual Style Prompting vs. StyleAligned
We additionally tested our score distillation on a different training-free style transfer method using diffusion model. Particularly, we adopt StyleAligned [22] instead of visual style prompting [30]. We found that our stylized score distillation works well with StyleAligned, and we did not find significant differences in the final results. The biggest difference occurs for harder cases mentioned in Section A.2 where StyleAligned produced results with more focus on surroundings of generated object. The comparison can be found in Figure 11.
Figure 11: Visual style prompting vs. StyleAligned.
C Derivation
Here we provide a detailed derivation of our stylized score distillation. Recall that our stylized score distillation aims to minimize $KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))$, which is equivalent to
$\min_{\theta}\mathbb{E}_{\varepsilon}$ $\left[\log(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta)))\right.$
$\left.-(1-\lambda)\log(p_{\phi}(\mathbf{z}_{t}\mid y))-\lambda\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right].$ (15)
@@ C Derivation
$\nabla_{\theta}KL$ $(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))$
$=\mathbb{E}_{\varepsilon}\left[\underbrace{\nabla_{\theta}\log(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta)))}_{(A)}\right.$
$\left.-(1-\lambda)\underbrace{\nabla_{\theta}\log(p_{\phi}(\mathbf{z}_{t}\mid y))}_{(B)}\right.$
$\left.-\lambda\underbrace{\nabla_{\theta}\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))}_{(C)}\right].$ (16)
@@
$\nabla_{\theta}\log(p_{\phi}(\mathbf{z}_{t}\mid y))=-\frac{\alpha_{t}}{\sigma_{t}}{\varepsilon_{\phi}}(\mathbf{z}_{t}\mid y)\frac{\partial{\mathbf{x}}}{\partial\theta},$ (17)
@@
$\nabla_{\theta}\log(\hat{p}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))=-\frac{\alpha_{t}}{\sigma_{t}}\hat{\varepsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})\frac{\partial{\mathbf{x}}}{\partial\theta}.$ (18)
@@
$\nabla_{\theta}\log q(\mathbf{z}_{t}|\mathbf{x})$ $=\left(\underbrace{\frac{\partial\log q(\mathbf{z}_{t}|\mathbf{x})}{\partial x}}_{\text{parameter score}}+\underbrace{\frac{\partial\log q(\mathbf{z}_{t}|\mathbf{x})}{\partial\mathbf{z}_{t}}\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{x}}}_{\text{path derivative}}\right)\alpha_{t}\frac{\partial\mathbf{x}}{\partial\theta}$
$=\left(\frac{\alpha_{t}}{\sigma_{t}}{\epsilon}-\frac{\alpha_{t}}{\sigma_{t}}{\epsilon}\right)\alpha_{t}\frac{\partial\mathbf{x}}{\partial\theta}=0.$ (19)
Similarly to DreamFusion [59], based on Sticking-the-Landing [60], we can discard the parameter score term and keep only the path derivative term.
@@
$\nabla_{\theta}\mathcal{L}_{SSD}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
$=\mathbb{E}_{t,\mathbf{z}_{t}\mid\mathbf{x}}\left[\omega(t)\frac{\sigma_{t}}{\alpha_{t}}\nabla_{\theta}KL(q(\mathbf{z}_{t}\mid\mathbf{x}=g(\theta))\parallel p_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s}))\right]$
$=\mathbb{E}_{t,\varepsilon}\left[\omega(t)\left((1-\lambda)\varepsilon_{\phi}(\mathbf{z}_{t}\mid y)+\lambda\hat{\varepsilon}_{\phi}(\mathbf{z}_{t}\mid y,\mathbf{s})-\varepsilon\right)\frac{\partial\mathbf{x}}{\partial\theta}\right].$ (20)
D GPT-4 Evaluation Template
We extend the GPTEval3D toolbox [74] to evaluate the style alignment between the generated 3D objects and the style reference image. The toolbox prompts the GPT-4 language model from OpenAI to perform the evaluation. The entire text prompt to GPT-4 is listed below. Instruction #6 is for style alignment evaluation. Figure 12 shows an example of the image grid sent to GPT-4 for evaluation. In Table 1, we present the detailed Elo scores obtained from this evaluation.
@@ 7. Overall, xxxxxx
Figure 12: An image grid sent to GPT for evaluation. The first and second column show the color rendering and normal map of the first method, and the third and fourth column are for the second method. Each row shows a camera view of the objects, and the last row shows the style reference. The prompt is ”a toy car” and the style image represents fire on a black background. GPT is asked to pick the better result out of the two presented methods.
@@
Table 1: Detailed results from GPTEval3D evaluation.